OpenStudy (anonymous):

how do i find the quotient of -88/-8?

OpenStudy (anonymous):

negative/negative=positive, so the answer will be positive

OpenStudy (texaschic101):

You divide -88 by -8 ,keeping in mind that when a negative is divided by a negative, it equals a positive. Example : 60/5 = 12 -60/5 = - 12 -60/-5 = 12 If the signs are the same, weather it is negative or positive, the result will be positive. If the signs are different, then the result will be negative.

OpenStudy (anonymous):

I think it's positive 11